The patient is an 8 year old child who weighs 55 lb. and is being treated...

  1. The patient is an 8 year old child who weighs 55 lb. and is being treated for streptococcal pharyngitis.

Ordered: Cephalexin Susp 200 mg PO q6h.

On hand: Cephalexin Oral Suspension USP 250 mg / 5 mL

Package Insert:

Pediatric patients: The usual recommended daily dosage for pediatric patients is 25 to 350 mg/kg in divided doses. For streptococcal pharyngitis in patients over 1 year of age and for skin and skin structure infections, to total daily dose may be divided every 12 hours.

A. Is the dosage ordered safe?   Show calculations that proves your choice.

  1. How will the medication be given?___________
  2. How often will the medication be given?_____________
  3. How much in mL of the medication to be given at prescribed time?________________
  4. What do you do if the dose is not a safe dose?_________________________________
  5. The patient is an 8 year old child who weighs 55 lb. and is being treated for streptococcal pharyngitis
Answer #1

ANSWER.

* Calculation of safe ordered dose.

        Ordered dose = 200 mg q6hr.

    Recommended daily dose = 25 to 350 mg/kg.

        Weight of the patient = 55 lb.

     Convert lb to kg for calculation

              1lb = 0.45 kg.

    So

      Weight of the patient = 55 * 0.45

                                       = 24.9 kg.

    Safe dose for a child weight 24.9 kg = 25 * 24.9 to 350 * 24.9

                                                               = 622.5 mg to 8715 mg in divided doses.

         Ordered dose = 200 mg q6hrly (4 times a day).

                                = 200 * 4

                                = 800mg / day.

* So the ordered dose is a safe dose for this children to treat streptococcal

      pharyngitis.

A) Cephalaxin used for this child is a oral suspension so it is given through oral

     route.

B) Physicians order is Cephalaxin q6hr,so it is given as 4 times a day.

C) Calculation of amount (ml) of drug given.

          Formula = Desired dose /Dose in hand * Quantity

     Here

            Desired dose (D) = 200 mg.

           Dose in hand (H) = 250 mg.

              Quantity (Q) = 5 ml.

As per above formula

       Amount(ml) of drug to be given = D/H * Q

                                                          = 200/250 * 5

                                                          = 0.8 * 5

                                                          =4 ml q6hr.

D) If the ordered dose is not safe withhold the dose and inform the physician to change

     the dose.